OpenStudy (anonymous):

There were 80 people in the parade. A total of 5/8 of the people in the parade played a musical instrument. How many people in the parade played a musical instrument?

OpenStudy (anonymous):

legendary saddist i need help with 3 more problems

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 }*\frac{ 80 }{ 1 }\to \frac{ 5*80 }{ 8*1 }\to \frac{ 400 }{ 8 }=50\]So fifty people

OpenStudy (anonymous):

so the answer is 50

OpenStudy (anonymous):

Yes, the answer is 50

OpenStudy (anonymous):

okay i have 1 more can u help me

OpenStudy (anonymous):

can you help me now?

OpenStudy (anonymous):

Dan had a roll of wrapping paper. He cut the paper from the roll into 15 equal pieces. Each piece was 2/3 yard long. What was the original length of the wrapping paper on the roll? A. 2 yards B. 5 yards C. 10 yards D. 15 yards

OpenStudy (anonymous):

me first

OpenStudy (anonymous):

this is my question tag someone ur self

OpenStudy (anonymous):

i asked first

OpenStudy (anonymous):

this is my question just go tag someone first

OpenStudy (anonymous):

i did it's legendarysadist

OpenStudy (anonymous):

@LegendarySadist canu help me with 1more

OpenStudy (anonymous):

@LegendarySadist help me first

OpenStudy (anonymous):

i am reporting u ur rude

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

@LegendarySadist canu help me

OpenStudy (anonymous):

Dan had a roll of wrapping paper. He cut the paper from the roll into 15 equal pieces. Each piece was 2/3 yard long. What was the original length of the wrapping paper on the roll? A. 2 yards B. 5 yards C. 10 yards D. 15 yards

OpenStudy (anonymous):

Yes yes i'm here. So it's\[\frac{ 15 }{ 1 }*\frac{ 2 }{ 3}\to \frac{ 15*2 }{ 1*3 }\to \frac{ 30 }{ 3 }=10\]So the answer is 10. Do you understand how I got it?

OpenStudy (anonymous):

yes do u wnat to hlp me with 1 more

OpenStudy (anonymous):

Sure, no problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Maria said that 3 × 4 = 12. Which equation correctly multiplies only prime numbers and equals the product of 12? A. 2 × 2 × 3 = 12 B. 2 × 2 × 2 × 3 = 12 C. 2 × 6 = 12 D. 6 × 6 = 12

OpenStudy (anonymous):

That would be A. Would you like me to explain why?

OpenStudy (anonymous):

no i got it but dont get mad i have 2 more

OpenStudy (anonymous):

do u knwo this

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

Yes, the questions just worded a little weird.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

But I believe it is A

OpenStudy (anonymous):

ok i have 1 more

OpenStudy (anonymous):

Ok, shoot

OpenStudy (anonymous):

First take the full integers and add them together. 1+4=?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

And now we add the fractions. since the common denominator is 100, we multiply 6/10 by 10/10.\[\frac{ 6 }{ 10 }*\frac{ 10 }{ 10 }\to \frac{ 60 }{ 100}\] Now that we have a common denominator, 60+7=?

OpenStudy (anonymous):

67

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

67

OpenStudy (anonymous):

Yes, I'm here. So the answer is \[5\frac{ 67 }{100}\]

OpenStudy (anonymous):

o hold on

OpenStudy (anonymous):

o hvae others can u help me i have 4 more

OpenStudy (anonymous):

Which steps could be used to solve this story problem? Krissy had 32 bagels. She gave an equal number of bagels to each child. There were 5 boys and 3 girls. How many bagels did the girls get altogether? A. Multiply 32 by 5. Multiply 32 by 3. Add the two products together. B. Add 5 and 3. Then multiply the sum by 32. C. Subtract 3 from 5. Then divide 32 by the difference. D. Add 5 and 3. Divide 32 by this sum. Then multiply the quotient by the number of girls.

OpenStudy (anonymous):

D for sure

OpenStudy (anonymous):

can you come back legendarysadist